3/4(2x+6)-3=3(x+3)
Answer by ethai18(3)

3/4(2x+6)-3=3(x+3)
3/2x+9/2-3=3x+9
3/2x+7.5=3x+9
-1.5x+7.5=9
-1.5x=2
x=-1.34